For a sample of size n-100, proportion p-0.6, and at a 95% confidence level, What is the upper bound of the proportion? Answer by stanbon(75887) 
 For a sample of size n-100, proportion p-0.6, and at a 95% confidence level, What is the upper bound of the proportion?
----------
Upper limit E= p-hat + 1.96*sqrt(pq/n)
ul = 0.6 + 1.96sqrt(0.6*0.4/100)
ul = 0.6 + 0.09602
ul = 0.69602
==========
Cheers,
Stan H.